PROGRESS TOWARDS CHIRPED-PULSE FOURIER TRANSFORM THZ SPECTROSCOPY

\underline{KEVIN O. DOUGLASS}, DAVID F. PLUSQUELLIC, NATIONAL INSTITUTE OF STAN-DARDS AND TECHNOLOGY, OPTICAL TECHNOLOGY DIVISION, GAITHERSBURG, MD 20899; EYAL GERECHT, NATIONAL INSTITUTE OF STANDARDS AND TECHNOL-OGY,ELECTROMAGNETICS DIVISION, BOULDER, CO 80305.

New opportunities are provided by the development of higher power THz frequency multiplier sources, the development of a broadband Chirped-Pulse FTMW spectroscopy technique at microwave^{a} and mmWave frequencies, and recently demonstrated heterodyne hot electron bolometer detection technology in the THz frequency region with near quantum noise-limited performance and high spectral resolution.^{b} Combining these three technologies and extending the chirpedpulse technique to 0.85 THz enables a host of new applications. NIST is currently pursing applications as a point sensor for greenhouse gases, volatile organic compounds, and potentially human breath. The generation and detection of phase stable chirped pulses at 850 GHz will be demonstrated. A description of the experimental setup and preliminary data will be presented for nitrous oxide.

aG.G. Brown, B.C. Dian, K.O. Douglass, S.M. Geyer, S. Shipman and B.H. Pate, Rev.Sci.Instrum. 79 (2008) 053103.

bE. Gerecht, D. Gu, L. You, K.S. Yngvesson, IEEE TRANSACTIONS ON MICROWAVE THEORY AND TECHNIQUES. 56, (2008) 1083.